A slow train travelling at 25 km/h leaves a station 1/2 h before another train travelling at 40 km/h. If the trains travel in the same direction on parallel routes, when does the faster train pass the slower train? What would x represent and what would y represent?
Answer by stanbon(75887)

A slow train traveling at 25 km/h leaves a station 1/2 h before another train traveling at 40 km/h. If the trains travel in the same direction on parallel routes, when does the faster train pass the slower train? What would x represent and what would y represent?
---------------
Faster train DATA:
rate = 40 km/h ; time = x-1/2 hrs ; distance = 40(x-(1/2)) = 40x-20 km
-----
Slower train DATA:
rate = 25 km/h ; time = x hrs ; distance = 25x km
------------------
Equation:
distance = distance
40x-20 = 25x
Subtract 25x from both sides.
Add 20 to both sides.
15x = 20
x = 4/3 hour = 80 minutes (time for the slower train)
x-(1/2) = 4/3-(1/2) = (8-3)/6 = 5/6 = 50 minutes (time for the faster train)
==================
Cheers,
Stan H.
===========